EDITORIAL

Terry Goodman

The end of a school year is always a time of reflection and evaluation. Here at MJMS we continue to be encouraged by the response to the journal. The number of manuscripts submitted to MJMS continues to grow and we have been most impressed by the quality of these papers. Likewise, the number of persons submitting problems and/or solutions to problems has increased this year.

Our referees continue to be a very important part of MJMS. Their prompt and insightful reviews have been most appreciated. At the end of the Editorial we have provided a listing of names of those persons who have served as referees for MJMS. We have also included a Referee Application Form on page 91 and we encourage anyone interested in serving as a referee for MJMS to fill out the form and return it to us.

Finally, the number of subscribers to MJMS is growing. This has the obvious financial benefit of helping to pay for the cost of the journal. Beyond that, it says that interest in MJMS is increasing and that ideas and issues are being considered by an ever broadening group of readers. We ask you to share MJMS with colleagues and friends and to encourage their participation.

We wish all of you a relaxing, refreshing, and productive summer.

List of Referees

Charles Allard, Polo R-VII High School

Mohammad Azarian, University of Evansville John Bernard, University of Texas–Pan American Sherralyn Craven, Central Missouri State University Albert Dixon, School of the Ozarks John Edgell, Jr., Southwest Texas State University Larry Eifler, University of Missouri–Kansas City Ekemezie Emeka, Columbia College Russell Euler, Northwest Missouri State University Joe Flowers, Northeast Missouri State University Ramesh Garimella, Northwest Missouri State University Victor Gummersheimer, Southeast Missouri State University Leon Hall, University of Missouri–Rolla Dennis Harmon, Missouri Southern State College Larry Johnson, Central Missouri State University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles Lamb, University of Texas–Austin William Livingston, Missouri Southern State College Gary McDonald, Northwest Missouri State University Rhonda McKee, Central Missouri State University Krzysztof Ostaszewski, University of Louisville James Petticrew, University of Texas–Pan American Robert Prielipp, University of Wisconsin–Oshkosh Don Redmond, Southern Illinois University Peter Shiue, University of Nevada–Las Vagas Shing So, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Alvin Tinsley, Central Missouri State University